# SWAT-NN: Simultaneous Weights and Architecture Training for Neural Networks in a Latent Space

### v2

@@ IV-2 Autoencoder Settings
We trained a multi-scale GPT-2 autoencoder with one encoder and four decoders for MLPs with 1-4 hidden layers. Each input MLP has two inputs, one output, and up to 7 neurons per hidden layer. Weights and biases are sampled from $[-5,5]$ and $[-1,1]$, respectively. For each MLP, 1000 inputs are uniformly sampled from $[-1,1]^{2}$, and the autoencoder is trained using the loss function in Section III-A5. Each training epoch includes 50,000 batches (batch size 64) and takes around 3 hours to complete on a single RTX A5000 GPU. In Section V-F, we also incorporate varying input and output dimensions by applying the same masking and padding strategy to the boundary layers.
@@ IV-3 Baselines
We train the DARTS model for 50 epochs to search for the architecture using bilevel optimization[11] procedure. ADMM is then applied with $\rho=2$ and a threshold of $10^{-1}$ to sparsify the weights. These hyperparameters are selected after increasing sparsity strength until performance on roughly one-third of the tasks dropped over 5% – our tolerable limit.
@@ IV-A Learning Sparse and Compact MLPs (Joint Optimization of Activation, Neuron Usage, and Sparsity)
In SWAT-NN setting, for sparsity penalty, we set $\lambda_{s}$ to $1\times 10^{-4}$, $\mu_{1}=0.1$ for the $\ell_{1}$ penalty and $\mu_{c}=0.01$ for the soft count term. For neuron compactness penalty, we use $\alpha=0.4$ and $\beta=0.001$. For activation function selection, we set $T_{\text{init}}$ = 1.0, $T_{\text{final}}$ = 0.01, $E_{\text{anneal}}$ = 3000. These hyperparameters are tuned using the last two benchmark functions (F53 and F54) from the CORNN dataset as a validation set. SWAT-NN training takes about 1 minute per task and per MLP configuration on a machine with a single RTX A5000 GPU, 128GB RAM, and a 56-core CPU.
Fig. 4 presents MSE vs non-zero weights for 28 benchmark functions, where each point represents a model configuration, and positions closer to the bottom-left indicate lower MSE and higher sparsity. SWAT-NN always identifies solutions closer to the optimal accuracy-sparsity trade-off, demonstrating its advantage in producing more compact and efficient networks.
Figure 4: MSE vs non-zero weights between DARTS+ADMM and SWAT-NN across 28 randomly selected benchmark functions. Each point represents a model configuration, plotted by its test MSE and number of non-zero weights.
Fig. 5 show the results on all dataset in terms of MSE and number of non-zero weights. For each function, we apply a two-step selection criterion: we first identify architectures whose MSE is within 5% of the minimum, and among them select the one with the fewest non-zero weights. While SWAT-NN and DARTS followed by ADMM achieve similar test accuracy, in most cases, SWAT-NN consistently discovers models with significantly fewer non-zero weights.
Figure 5: Top: Test MSE of the best-performing models identified by DARTS+ADMM and SWAT-NN across all 54 functions. Bottom: Corresponding number of non-zero weights. Function labels are color-coded to indicate which method yields a more sparse and compact model.
@@ IV-B Search for Optimal MLPs with Activation Functions
Fig. 7 further compares our method to the traditional training baseline described in Section IV-3. Our method achieves performance comparable to MLPs trained using traditional training method with fixed activation functions in most cases. While the accuracy is similar, the ability to flexibly assign activation functions per neuron enables our method to produce more compact and sparse networks, as demonstrated in Section IV-A and Section IV-C.
@@ V-B Comparison Between CNN-Based and GPT-Based Autoencoders
A prior work [17] used a CNN-based autoencoder. In this work, we use a GPT-based autoencoder, which we find more effective at capturing functional representations. Both autoencoders are trained under the same basic setting – without bias terms, neuron masking, or activation function selection (all neurons use sigmoid). Using the same training configuration, we compare the best MSE achieved in downstream MLP search on CORNN dataset. As shown in Fig. 11, the GPT-based model consistently outperforms the CNN-based one.
Figure 11: Comparison of minimum test MSE across benchmark functions using CNN-based and GPT-based autoencoders.
@@ V-C Functional Approximation by the Autoencoder
We hypothesize that this is because the embedding space only needs to provide a sufficiently smooth and semantically organized landscape, such that gradient descent can navigate toward better-performing regions. The autoencoder does not need to achieve perfect reconstruction, but rather serve as a functional embedding mechanism. Improving the fidelity of the autoencoder would further enhance the smoothness and precision of the search space, leading to improved performance in the downstream optimization.
@@ V-D Different hyperparameter choices
Small Penalty $1\times 10^{-5}$ $1\times 10^{-1}$ $1\times 10^{-4}$
@@
Figure 13: Test MSE (top) and number of non-zero weights (bottom) for the best-performing networks under three different levels of sparsity and compactness regularization.
@@ V-E Change in size of the embedding space
To assess whether a large embedding dimensionality is necessary, we train autoencoders to embed 1–4 hidden layer MLPs with fixed structure: no bias terms, no activation functions, and exactly 5 neurons per layer. Each autoencoder is trained for 50 epochs, with 640,000 sampled MLPs per epoch. Fig. 14 compares performance on all regression tasks across different embedding sizes. In nearly all tasks, the $5\times 768$ embedding yields lower or equal MSE loss compared to smaller alternatives ($5\times 256$ and $1\times 512$). This suggests that a sufficiently large embedding space is beneficial for capturing functional variations in MLPs. Accordingly, we use a $7\times 768$ embedding in our main experiments.
Figure 14: Test MSE for the best MLPs found using embedding vectors of different dimensions on the CORNN dataset. Each curve corresponds to a specific embedding size: $5\times 768$, $5\times 256$, and $1\times 512$.
@@ VI Conclusion
We presented SWAT-NN, a novel framework that simultaneously optimizes both neural architectures and their weights within a universal and continuous embedding space. Through extensive experiments on all 54 regression tasks in the CORNN dataset, we demonstrated that SWAT-NN consistently yields performant networks with desirable sparsity and compactness, achieving comparable or better accuracy while producing more compact and sparse neural networks than state-of-the-art methods.

### v3

@@ IV-2 Autoencoder Settings
We trained a multi-scale GPT-2 autoencoder with one encoder and four decoders for MLPs with 1-4 hidden layers. Each input MLP has two inputs, one output, and up to 7 neurons per hidden layer. Weights and biases are sampled from [5,5][-5,5] and [1,1][-1,1], respectively. For each MLP, 1000 inputs are uniformly sampled from $1,1]^{2}$, and the autoencoder is trained using the loss function in Section III-A5. Each training epoch includes 50,000 batches (batch size 64) and takes around 3 hours to complete on a single RTX A5000 GPU. In Section V-F, we also incorporate varying input and output dimensions by applying the same masking and padding strategy to the boundary layers.
@@ IV-3 Baselines
We train the DARTS model for 50 epochs to search for the architecture using bilevel optimization[liu2018darts] procedure. ADMM is then applied with $\rho=2$ and a threshold of $10^{-1}$ to sparsify the weights. These hyperparameters are selected after increasing sparsity strength until performance on roughly one-third of the tasks dropped over 5% – our tolerable limit.
@@ IV-A Learning Sparse and Compact MLPs (Joint Optimization of Activation, Neuron Usage, and Sparsity)
In SWAT-NN setting, for sparsity penalty, we set $\lambda_{s}$ to $1\times 10^{-4}$, $\mu_{1}=0.1$ for the $\ell_{1}$ penalty and $\mu_{c}=0.01$ for the soft count term. For neuron compactness penalty, we use $\alpha=0.4$ and $\beta=0.001$. For activation function selection, we set $T_{\text{init}}$ = 1.0, $T_{\text{final}}$ = 0.01, $E_{\text{anneal}}$ = 3000. These hyperparameters are tuned using the last two benchmark functions (F53 and F54) from the CORNN benchmark suite as a validation set. SWAT-NN training takes about 1 minute per task and per MLP configuration on a machine with a single RTX A5000 GPU, 128GB RAM, and a 56-core CPU.
Fig. 4 presents MSE vs non-zero weights for 20 datasets, where each point represents a model configuration, and positions closer to the bottom-left indicate lower MSE and higher sparsity. SWAT-NN always identifies solutions closer to the optimal accuracy-sparsity trade-off, demonstrating its advantage in producing more compact and efficient networks.
Figure 4: MSE vs non-zero weights between DARTS+ADMM and SWAT-NN across 20 randomly selected datasets. Each point represents a model configuration, plotted by its test MSE and number of non-zero weights.
Fig. 5 show the results on all datasets in terms of MSE and number of non-zero weights. For each dataset, we apply a two-step selection criterion: we first identify architectures whose MSE is within 5% of the minimum, and among them select the one with the fewest non-zero weights. While SWAT-NN and DARTS followed by ADMM achieve similar test accuracy, in most cases, SWAT-NN consistently discovers models with significantly fewer non-zero weights.
Figure 5: Top: Test MSE of the best-performing models identified by DARTS+ADMM and SWAT-NN across all 54 datasets. Bottom: Corresponding number of non-zero weights. Function labels are color-coded to indicate which method yields a more sparse and compact model.
@@ IV-B Search for Optimal MLPs with Activation Functions
Fig. 7 further compares our method to the traditional training baseline described in Section IV-3. Our method achieves performance comparable to MLPs trained using traditional training method with fixed activation functions in most cases. While the accuracy is similar or slightly lower, the ability to flexibly assign activation functions per neuron enables our method to produce significantly more compact and sparse networks, as demonstrated in Section IV-A and Section IV-C.
@@ V-B Comparison Between CNN-Based and GPT-Based Autoencoders
As an ablation, we compare CNN-based and GPT-based autoencoders to examine which is more effective at capturing functional representations. Both autoencoders are trained under the same basic setting – without bias terms, neuron masking, or activation function selection (all neurons use sigmoid). Using the same training configuration, we compare the best MSE achieved in downstream MLP search on CORNN benchmark suite. As shown in Fig. 11, the GPT-based model consistently outperforms the CNN-based one.
Figure 11: Comparison of minimum test MSE across benchmark function datasets using CNN-based and GPT-based autoencoders.
@@ V-C Functional Approximation by the Autoencoder
We hypothesize that the embedding space only needs to provide a sufficiently smooth and semantically organized landscape, such that gradient descent can navigate toward better-performing regions. It does not need to achieve perfect reconstruction, but rather serve as a functional embedding mechanism. Improving the fidelity of the autoencoder would enhance the smoothness and precision of the search space, leading to better performance in the downstream optimization.
@@ V-D Different hyperparameter choices
Small Penalty $\times 10^{-5} 1×1$$\times 10^{-1}$ $1\times 10^{-4}$
@@
Figure 13: Test MSE (top) and non-zero weights (bottom) for the best networks under three sparsity and compactness regularization levels.
@@ V-E Change in size of the embedding space
To assess whether a large embedding dimensionality is necessary, we train autoencoders to embed 1–4 hidden layer MLPs with fixed structure: no bias terms, no activation functions, and exactly 5 neurons per layer. Each autoencoder is trained for 50 epochs, with 640,000 sampled MLPs per epoch. Fig. 14 compares performance on all regression datasets across different embedding sizes. In nearly all datasets, the $5\times 768$ embedding yields lower or equal MSE loss compared to smaller alternatives ($5\times 256$ and $1\times 512$). This suggests that a sufficiently large embedding space is beneficial for capturing functional variations in MLPs. Accordingly, we use a $7\times 768$ embedding in our main experiments.
Figure 14: Test MSE for the best MLPs found using embedding vectors of different dimensions on the CORNN benchmark suite. Each curve corresponds to a specific embedding size: $5\times 768$, $5\times 256$, and $1\times 512$.
@@ VI Conclusion
We presented SWAT-NN, a novel framework that simultaneously optimizes both neural architectures and their weights within a universal and continuous embedding space. Through extensive experiments on all 54 regression datasets in the CORNN benchmark suite, we demonstrated that SWAT-NN consistently yields performant networks with desirable sparsity and compactness, achieving comparable or better accuracy while producing significantly more compact and sparse neural networks than state-of-the-art methods.